1. The sugar jar has 4 cups of sugar when full. Joe ate 1% of the sugar in the jar. Explain what this means. Is this possible? Why or why not
2. On a recent test Ron only got 2 questions correct out of 10. The teacher let him retake it and he got 350% more correct. Explain what this means in your own words. Is this possible?
3. A cell phone commercial stated that 110% of your calls will be cheaper with their company. Explain what this means in your own words. Is this possible?
4. Bleu read 0.1% of the books at her school library. Explain what this means in your own words. Is this possible? Why or why not?

Answer by macston(5194)

.
1 percent means 1/100
1/100(4cups)=4/100 cup
(4/100 cup)(236.59g/cup)=9.46g
There are 39g sugar in a 12 ounce coke, so Joe could easily consume 9.46 grams.